OpenStudy (anonymous):

Simplify the given expression. square root of negative 64 over open parentheses 7 minus 6 I close parentheses minus open parentheses 2 minus 2 i

OpenStudy (anonymous):

words=numbers please :(

OpenStudy (anonymous):

ill write it out now

OpenStudy (anonymous):

\[\sqrt{-64\div} \left( 7 - 6i \right) - \left( 2- 2i \right)\]

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt{-64}}{(7-6i)}-(2-2i)\] do you mean this ?

OpenStudy (anonymous):

@cuddlepony there's one "i" after number 6

OpenStudy (anonymous):

no its the sqrt of -64 divide by everything else

OpenStudy (anonymous):

ahhh ... sweet . Now I see ;) \[\huge \frac{\sqrt{-64}}{(7-6i)-(2-2i)}\] this must be then ;) ... let me try ;)

OpenStudy (anonymous):

AAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAAWWWWWWWWWWWWWWWWWWW :@:@:@:@:@:@ F#C#K@NG GOOGLE CRASHED :@ EVERYTHING LOST !! ....

OpenStudy (anonymous):

I'll write again .. hold on

OpenStudy (anonymous):

oh no :((

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt{-64}}{(7-6i)-(2-2i)}=\frac{i\sqrt{2\cdot 2\cdot 2}}{(7-6i)+(2i-2)}\] I'll post it by parts... so it would not be lost.. :S

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

it is ..\[\LARGE i \sqrt{4\cdot 4\cdot 4}\]

OpenStudy (anonymous):

I'll do it faster I have an idea !!! ... wait.

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

http://www.twiddla.com/806371

OpenStudy (anonymous):

do you see it? .. ... do you have to find imaginary part and real part or only to simplify ? :) LATEX SUCKS ! lol

OpenStudy (anonymous):

@mscrosscountry still there?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

Glad to help :)